OpenStudy (dallaslifebaby01):

A car uses one gallon of gasoline for every 20 miles it travels. If a gallon of gasoline costs $3.98, how much will the gas cost, to the nearest dollar, to travel 180 miles?

OpenStudy (dallaslifebaby01):

i chose 36 the answer choices are 9 36 45 80

OpenStudy (dallaslifebaby01):

@hara3104 @FaiqRaees @Jaynator495 @sammixboo random tags

OpenStudy (imstuck):

Set this up in a ratio situation, or you could call it an equality situation, or proportions. Either way, it involves fractions that are equivalent to one another. Set it up like this: \[\frac{ $3.98 }{ 20 }=\frac{ x }{ 180 }\]

OpenStudy (imstuck):

Money goes on the top and miles on the bottom. We know that it takes $3.98 to travel 20 miles, we need to find out how much money, x, it takes to go 180 miles. Make sure the money is on top together (or bottom..as long as you are consistent) and the miles on the other. Cross multiply to solve for x.

OpenStudy (imstuck):

$3.98 * 180 = 20 * x

OpenStudy (imstuck):

Rounding up you do get 36, so good job!